Grade 11General Physics

Time period of a pendulum depends on l, g and m. Find the formula for time period

Lasdfggg
9 Years agoGrade 11

1 Answer

Vikas TU
9 Years ago
 
Dear Student,
Pendulum takes after SHM. 
At the point when the pendulum has swung and makes an edge  with the underlying vertical harmony position, 
Frestoring = matangential 
mg sinθ = matangential 
The curve length, s, is specifically corresponding to the extent of the focal point, θ, 
s = rθ. 
=> s = Lθ, where θ must be measured in radians. 
Substituting into the condition for SHM 
Frestoring= - ks 
mg sinθ = - k(Lθ) 
Settling for the "spring steady" or k for a pendulum yields 
mg sinθ = k(Lθ) 
k = mg sinθ/Lθ 
(lim)┬(n→infinity)⁡〖sinx/x〗 = 1 
k=mg/L 
Substituting k in the SHM condition for day and age, 
T=2π√(m/k) 
=>T=2π√(mL/mg) 
=>T=2π√(L/g).
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)